OpenStudy (anonymous):

Explain the meaning of absolute value of a complex number. What is the absolute value of –2 + 3i?

OpenStudy (anonymous):

The absolute value of a complex number is that numbers distance from zero on a corrdinate axis...

OpenStudy (anonymous):

Okay that makes sence, how do you find it?

OpenStudy (anonymous):

using the pythagorean theorem you have sqrt( -2^2 + 3^2 ) = abs(-2+3i)

OpenStudy (anonymous):

the coordinate plane has a real component and an imaginary component, corresponding, respectivly, to the x and y axis of the more conventional version

OpenStudy (anonymous):

-2 is your "x" value and 3 is your "y" value, other wise this is a pythagorean theorem problem with a right triangle

OpenStudy (anonymous):

so the absolute value is 1?

OpenStudy (anonymous):

sqrt(-2^2 + 3^2) = sqrt(13)

OpenStudy (anonymous):

okay im confused now

OpenStudy (anonymous):

how did you get the 13

OpenStudy (anonymous):

-2^2 = 4 and 3^2 = 9, so -2^2+3^2 = 13